a pharmascist has a 40% acid solution and a 25% acid solution. How many liters of each must be mixed to form 147 liters of a 34% acid solution?
Answer by stanbon(75887)

A pharmacist has a 40% acid solution and a 25% acid solution.
How many liters of each must be mixed to form 147 liters of
a 34% acid solution?
------------------------
Equation:
acid + acid = acid
0.40x + 0.25(147-x) = 0.34*147
----
Multiply thru by 100 to get:
40x + 25*147 - 25x = 34*147
----
15x = 9*147
x = (3/5)(147)
x = 88.2 liters (amt. of 40% solution needed)
147-x = 58.8 liters (amt. of 25% solution needed)
====================================================
Cheers,
Stan H.
=====================